Three forces F1, F2 and F3 acting at O are in equilibruim, and the angle at inclination is 60, If the magnitude of F1 is 100N and F2 is 5N. calculate for F3


1
Expert's answer
2021-03-09T15:27:13-0500

if the force "F_1=100\ N" and is directed at an angle of "60\u00b0" to the the positive x-axis, "F_2=5\ N"


along the negative Y-axis.


"F_x=100\cdot\cos60°=50\ (N)"


"F_y=100\cdot\sin60°-5=81.6\ (N)"


"F_3=\sqrt{50^2+81.6^2}=95.7\ (N)"


"\alpha=\tan^{-1}\frac{81.6}{50}=58.5°"


The force "F_3=95.7\ N" and is directed at an angle of "238.5\u00b0" to the the positive x-axis. Answer